# Shahriar Noroozizadeh

### v2

@@ C.3 Tiny graphs
C.4 Additional experiments on path-star geometry
In the main paper, we present heatmaps showcasing the distance between the leaf nodes and the first node in every path. We consolidate these below and also provide an additional path-to-path distance heatmap.
Leaf–first-hop distance (Fig.˜25). In Fig.˜25 (an extended version of the heatmaps in Fig.˜6), entry $(i,j)$ is the cosine distance between the leaf embedding of path $i$ and the first-hop (indecipherable token) embedding of path $j$. Diagonal entries are low (a leaf lies close to its correct first hop), while off-diagonals are higher. This structure explains why the first-token-only objective succeeds in-weights (Fig.˜4-(Right)): the $k$-fold composition map reduces to a local geometric step in the learned representation.
Path–by–path distance (Fig.˜26). Instead of only analyzing the distance between the leaf and first nodes, we analyze the distance across all pairs of nodes in a path. For each pair of paths $(i,j)$, we compute the mean distance between all node embeddings on path $i$ and all node embeddings on path $j$.
  • Diagonal — Intra-Path Distance ($i=j$): This value is the average distance between all unique pairs of distinct nodes within a single path. It measures how tightly clustered the path’s nodes are. A smaller value indicates higher cohesion.
    $D_{i,i}=\frac{1}{\binom{\ell}{2}}\sum_{1\leq k<m<\ell}d(v_{k}^{(i)},v_{m}^{(i)})$
  • Off-diagonal — Inter-Path Distance ($i\neq j$): This value is the average distance between all nodes of one path and all nodes of another. It measures how separated two distinct paths are. A larger value indicates greater separation.
    $D_{i,j}=\frac{1}{\ell^{2}}\sum_{k=1}^{\ell-1}\sum_{m=1}^{\ell-1}d(v_{k}^{(i)},v_{m}^{(j)})$
We find a similar clustering of nodes within paths here, although the diagonal is generally less vivid. For the model trained only on edge-memorization, we do not see the diagonal at all.
(a) Edge-Memorized and
Full-Path-Trained
(b) Edge-Memorized and
Hardest-Token-Trained
(c) Only Edge-Memorized
Figure 25: Evidence of global geometry in path-star task: Leaf-first-token cosine distance between node embeddings. We present again the heatmaps from Fig.˜6 with an additional heatmap in the middle. Entry $(i,j)$ is the mean cosine distance between the leaf token in (an unseen) path $i$ (row) and first/hardest token on (an unseen) path $j$ (col). Each heatmap corresponds to a different training objective: Left: trained on edges and path-finding task $(\mathcal{D}_{\texttt{edge}}\cup\mathcal{D}_{\texttt{path}}^{\to})$; Middle: trained on edges and hardest-token-finding task (not presented in the main paper); Right: edges only $(\mathcal{D}_{\texttt{edge}})$. We find that even on these unseen paths, the leaf and first token embeddings cluster together, regardless of whether path-finding supervision exists; however, the geometry is strongest with global supervision (however, in Mamba SSM, even the locally supervised model shows strong geometry; see Fig.˜15).
(a) Edge-Memorized and
Full-Path-Trained
(b) Edge-Memorized and
Hardest-Token-Trained
(c) Only Edge-Memorized
Figure 26: Evidence of global geometry in path-star task: Pathwise average cosine{cosine} distance between node embeddings. While in Fig.˜25, we reported the cosine distance between the embeddings of the terminal nodes of a pair of (unseen) paths, here we consider an average over all nodes in those (unseen) paths. In particular, entry $(i,j)$ is the mean cosinecosine distance between nodes on path $i$ (row) and nodes on path $j$ (col). On the first two settings, as before, we find that the diagonal cells are low, implying closer embeddings. Thus, a geometry has emerged on unseen paths. However, unlike in Fig.˜6, we do not see any such signal when trained only on the edge memorization task.
UMAP projection (Fig.˜27, zoomed in version of Fig.˜6, middle). A different way to establish geometry is to directly visualize the paths. As done in Fig.˜6, we do this by projecting the token embeddings of all nodes with default UMAP settings (neighbors=15, min_dist=0.1). A zoomed in version of this is presented in Fig.˜27. We exclude the root token embedding in these projections since that is common to all paths. We find that different paths form well-separated clusters, and within each path cluster, nodes tend to arrange from leaf toward ${v}_{\tt{root}}$.
Figure 27: Zoomed in version of Fig.˜6: UMAP projection of token embeddings exhibits path-star topology. Each point is a node embedding; color indicates path identity. Different paths form separated clusters; the central node ${v}_{\tt{root}}$ is excluded since it is shared across all paths. The graph has degree $10^{4}$ and path length 66 (which includes the root). Axes are UMAP components (arbitrary units). Note that we have re-used the same color for multiple paths.
@@ Appendix D Edge supervision and training dynamics
There are tangential aspects of our training that are worth elaborating on: the roles of reverse edges (§D.1), of pause tokens (§D.2), of interleaving edge-memorization (§D.3), of weight-tying (§D.4), and of the learning order of tokens (§D.5).
@@ D.1 The role of reverse edges
Reverse edges seem to play a nuanced role in our observations. On the large path-star task in §2, we find it necessary to augment training on the reverse edges; on the other hand, for the tiny graphs, a geometry arises even without these reverse edges. Perhaps, reverse edges are needed for larger tasks; or perhaps, they are necessary to perform implicit reasoning and retrieval. We leave it for future work to gain greater clarity on this effect, which is tied to the reversal curse [18, 9].
@@ D.1.1 The critical role of reverse edges in the large path-star task
(Role of reverse edges) We find in Fig.˜28 that:
@@
    A Transformer trained on only the forward edges, struggles on both forward and reverse path-finding tasks (see the middle color in Fig.˜28).
@@
    A Transformer trained on only the reverse edges, achieves non-trivial accuracy on the reverse path-finding task; however, it fails on the forward path-finding task (see the third color in Fig.˜28).
@@
We note that the asymmetry between forward and reversed path-generation tasks stems from their algorithmic complexity. The reversed path generation is algorithmically trivial on path-star graphs because each node has a unique predecessor along any target path—the model simply needs to follow the unique backward edges. Forward generation, however, requires planning (examine each outgoing path) or lookahead (track the reverse path without explicit chain-of-thought and reverse it). Indeed, for the in-context task of B&N’24, the model fails on the forward task, but strikingly succeeds on the reverse task, as corroborated in Fig.˜29 (right). Even in the in-weights setting Fig.˜29 (left), the reverse path is generally quicker to learn and yields higher accuracy.
@@
D.1.2 Tiny graphs with uni-directional edge-memorization
In Fig.˜30, we revisit our tiny graphs in §C.3, and examine the embeddings of a Transformer when it memorizes only one direction of the edges. We find that a geometry still arises, although a bit weaker.
Figure 30: Embeddings of a Transformer with bi-directional vs. uni-directional edge memorization. With our smaller graphs, we find a geometry arise regardless of whether the model is made to memorize both or only one direction of each edge. However, the geometry is weaker (e.g., for the grid graph) under uni-directional memorization.
@@ D.2 Pause tokens for computational slack
In the same in-weights path-finding task of §2, we find that it is helpful to insert pause tokens [25, 57] to achieve quicker accuracy gains during training. Pause tokens are added by appending dummy tokens to the prefix of the path-finding task both during training and inference.
Fig.˜31 shows that adding a short sequence of pause tokens after the prompt reliably boosts exact-match accuracy across graphs, for a given amount of training time. Increasing the number of pause tokens increases speed of convergence.
@@
In all our experiments, we have interleaved edge-memorization examples with path-finding examples. An alternative training method would be a two-phased approach, where we first enforce edge-memorization, and then follow up by finetuning on the path-finding task. We found this to be less stable, e.g., the model achieves a peak accuracy momentarily, only to deteriorate dramatically right after. This is a manifestation of the well-known effect that finetuning has on parametric memory [95, 103]. Since this is a confounding effect, we do not choose this regime for our experiments.
However, we confirm that even in this regime, our models do achieve a high peak accuracy (see Fig.˜32). The fact that the composition task is learnable in this regime implies that the edge-pretrained model must have come with an adequate global geometry despite being trained only on local supervision (Section˜3.1).
@@
D.4 Effect of weight-untying
As discussed in §4.4, weight-untied models do not produce an apparent geometry or exhibit high node-to-node cosine-similarities, since it corresponds to SVD directions. We demonstrate this for the Transformers in Fig.˜11 and Fig.˜34 and Node2Vec in Fig.˜10 and Fig.˜33.
(a) Tiny Path-Star
(b) Tiny Grid
(c) Tiny Cycle
(d) Tiny Irregular
Figure 33: Node-node cosine similarity vs. adjacency matrix for weight-untied Node2Vec. As discussed in §4.4, observe that the node-to-node cosine-similarity matrix appears to be a bland near-zero matrix, lacking any information about the multi-hop connectivity of the underlying graph.
(a) Tiny Path-Star
(b) Tiny Grid
(c) Tiny Cycle
(d) Tiny Irregular
Figure 34: Node-node cosine similarity vs. adjacency matrix for weight-untied Transformer. As discussed in §4.4, observe that the node-to-node cosine-similarity matrix appears to be a bland near-zero matrix, lacking any information about the multi-hop connectivity of the underlying graph.
D.5 Learning order of tokens
For clarity, in Fig.˜35, we provide a side-by-side contrast between the order in which tokens are learned in the in-weights setup (with next-token prediction) and the in-context setup (with multi-token prediction).
(a) In-context (teacherless)
(b) In-weights
Figure 35: Learning dynamics per token. (a) In the in-context setting of $\mathcal{G}_{2,5}$ (trained with a multi-token, teacherless objective since standard next-token prediction fails), later tokens are learned first indicating strong reliance on future-token signals. (b) In the in-weights setting of $\mathcal{G}_{10^{4},6}$ with next-token prediction, token accuracies rise largely in tandem (or in a somewhat confusing order); the first token is not selectively driven by future targets.
@@ E.1 Empirical failure of composition learning under associative memory
As discussed in §2.3, it is well-known that certain compositional tasks are hard to learn empirically; theoretically, this has been proven in a certain sense. However, these prior discussions involve composing information available in the context, rather than in the weights. To test this intuition in our in-weights composition task, we design an experiment where we freeze the embeddings of a Transformer and train it on our path-star task. If the model succeeded in this task, it may mean one of two things: either the model develops a geometric memory in a subsequent layer, or the model does in fact efficiently learn how to compose associative matrix operations—going against our intuition derived from prior limits on composition learning. However, we find that even after $50,000$ training epochs (using the same GPT-mid architecture described in Table 2 and §B.2.1, except with frozen token embeddings; and the optimization hyperparameter grid search reported in §B.3), the model fails to learn the in-weights path-star task. We believe, this is preliminary evidence that associative memory indeed struggles to learn compositional in-weights tasks. A fleshed-out proof of this negative result is left for future work.
@@
In datasets where redundancies exist, the complexity of a lookup table scales quickly with the training set size (say $n$), whereas the more succinct solution does not (or at worst, grows polynomially slower). For example, if the data is linearly separable in some constant dimensionality, the linear classifier can be described in a constant number of bits, whereas a lookup table (such as a nearest neighbor model) would require $n$ bits. However, this wide disparity in complexity does not necessarily surface in our setting, which is a memorization task without redundancies. Concretely, at least in terms of bits and $\ell_{2}$ norms, there are simple graphs where both the geometric and associative views are equally complex—there is no factor of $n$, which in our case should be the edge count or the vertex count of our graph. We informally prove this below.
We first roughly derive the bit and norm complexity for a general graph, showing how an associative memory scales with the edge count, whereas a geometric memory scales with the vertex count. Then, we argue how this resolves to similar values for graphs like the path-star or a cycle, where the edge count and the vertex count are the same (almost).
@@
In the standard associative memory view that we have discussed, associations are represented through the function ${\bf{\Phi}}(v)^{T}{\bf{W}}_{\tt assoc}{\bf{\Phi}}(u)$. However, weight-tied dual encoder models like Node2Vec can only represent functions of the form ${\bf{\Phi}}(v)^{T}{\bf{\Phi}}(u)$. When weight-tied, this precludes the form of associative memory we care about; but it still allows two forms of associative memories under special conditions: if the graph is bipartite (Proposition˜3) or if the embedding dimensionality is sufficiently large (in which case, the construction is contrived; see Proposition˜5). We can also show that non-weight-tied models can represent any graph associatively (see Proposition˜4). In all these results below, we will show that our constructions achieve a logit of $1$ on adjacent vertices, but a logit of 0 on any non-adjacent vertices. In this sense, only local associations are captured, whereas no global geometry is.
@@
For each node, we take its corresponding row from the “incidence” matrix. In other words, the embedding is $|E|$-dimensional, where the $i$th dimension corresponds to the $i$th edge; for edge $(u,v)$, the embedding of $u$ and $v$ both are set to $1$ along the dimension corresponding to $(u,v)$. Notationally, if the edges are indexed as $1,2,\ldots$, then ${\bf{\Phi}}(u)_{i}=\mathbf{1}[u\in e_{i}]$, where $\bf{1}$ is the indicator function. Then, we have that ${\bf{\Phi}}(u)\cdot{\bf{\Phi}}(v)=\mathbf{1}[(u,v)\in E]$. Thus, the dot products capture only local information. ∎

### v3

@@ C.3 Tiny graphs
C.3.1 Effect of explicit regularizers and self-edges
We provide an analysis of how the geometry is affected by various aspects of the optimization setting: adding self-edges, varying the initialization, the learning rate, weight decay, and dropout. For all experiments in this section, we train for 10,00010{,}000 optimization steps to ensure comparability across settings. For this study, we use the TinyNN architecture with embedding dimension $512$ trained on the easy-to-visualize cycle graph. For each setting, we provide a visualization of (a) the embeddings, (b) the cosine heatmap of the embeddings, and (c) the embeddings’ projections onto the graph’s eigenvectors, along with the geometric margin (defined in §B.5), which when positive confirms that the storage is indeed geometric.
Our main observations are as follows:
  1. 1.
    Self-edges: Adding self-edges to the training graph makes the visualizations significantly cleaner and less zigzaggy (see Fig.˜25). This aligns with our understanding of the negative eigendirections in §4.4: adding self-edges makes such negative directions less dominant by bringing them closer to zero. Indeed, we see this in the eigenvector projections in Fig.˜25(c).
  2. 2.
    Initialization: The scale of initialization has a strong effect on what is learned by the model. Specifically, our default GPT-style initialization of embeddings ($\mathcal{N}(0,0.2^{2})$) consistently provided more geometric results, even without any explicit regularization (Fig.˜26). Whereas, a unit-normal initialization of the embeddings (Fig.˜28) resulted in a wider range of behaviors, more sensitive to training conditions e.g., in Fig.˜28, we show a setting that produces an associative storage.
  3. 3.
    Learning rate: We found that higher learning rate tends to improve the geometry. This result is most pronounced for the unit-normal initialization of embeddings with $\mathcal{N}(0,1)$, (compare Fig.˜28 with the higher learning rate Fig.˜29), but is also seen for GPT-style initialization (compare Fig.˜27 with the higher learning rate Fig.˜26).
  4. 4.
    Explicit regularization: On the unit-normal initialization of embeddings, which we found to be sensitive to optimizer hyperparameters, we find stronger geometries with increasing the weight decay (Fig.˜30); increasing dropout also helps, though the effect is less pronounced (Fig.˜31). As noted earlier, on the GPT-style initialization, such regularizers were not necessary (Fig.˜26).
  5. 5.
    Forward vs. reverse edges: The geometry does not seem contingent on there being both directions of the edges on the training set for our Transformer model (Fig.˜32)
Graph memorization as a minimal example of grokking. These observations point to the fact that geometric memorization of a graph dataset is a minimal instance of grokking, a phenomenon studied for more complex algorithmic datasets (where closed form solutions to the weights are hard to obtain). The helpful effect of weight decay may be related to its effects on grokking [Liu et al., 2023] or rank-minimization [Wang and Jacot, 2024, Galanti et al., 2025, Yunis et al., 2024]; similar rank-minimizing effects have been identified for dropout [Cavazza et al., 2018] and learning rate [Andriushchenko et al., 2023]. The effect of initialization has also been pointed out in grokking [Liu et al., 2023], and connected to the idea of lazy vs. rich regimes [Kumar et al., 2024] from the perspective of neural tangent kernels, where rich feature-learning is analogous to not geometric representations. However such connections between grokking and the lazy/rich dichotomy is contentious because associative memory, unlike lazy learning, requires the weights to move substantially; similar points have been made in [Zheng et al., 2025, Chou et al., 2025].
(a) PCA
(b) Emb. Heatmap vs. Adjacency
(c) Eigenvector Projections
Figure 25: Self-edge reduces zig-zagging: When self-edges are added to the training set, the model relies less on the bottommost eigendirections that are zig-zagging. This is evidenced by the smooth heatmap and the top-heavy barplot; compare this with zigzagging heatmaps and more bottom-heavy barplot in the no-self-edge setting of Fig.˜26. For this setup, we use a GPT-style initialized network with LR: 0.01; Embedding Init: GPT-Default ($\mathcal{N}(0,0.02^{2})$), Weight Decay: Off, Dropout: Off
(a) PCA
(b) Emb. Heatmap vs. Adjacency
(c) Eigenvector Projections
Figure 26: Geometry arises without regularization: Without self-edges or weight decay or dropout, we still see a global geometry in the heatmap, and a skew (towards the highest magnitude) eigendirections. Quantitatively, the geometric margin of this model is indeed positive ($7.87$), affirming that the graph is indeed stored in the embedding matrices. For this setup, we use a LR: 0.01; Embedding Init: GPT-Default ($\mathcal{N}(0,0.02^{2})$), Weight Decay: Off, Dropout: Off
(a) PCA
(b) Emb. Heatmap vs. Adjacency
(c) Eigenvector Projections
Figure 27: Lower learning rate weakens the global geometry: Reducing the learning rate from $10^{-2}$ in Fig.˜26 to $10^{-4}$ still results in a geometry, but the eigendirections are less skewed towards the global directions, and the heatmap weaker. For this setup, we use LR: 0.0001; Embedding Init: GPT-Default ($\mathcal{N}(0,0.02^{2})$), Weight Decay: Off, Dropout: Off
(a) PCA
(b) Emb. Heatmap vs. Adjacency
(c) Eigenvector Projections
Figure 28: Unit-Normal-init learns less geometric solution: Changing the initialization in Fig.˜27 results in what is now a purely-associative solution: the heatmap indicates orthogonal embeddings and equally distributed along all directions. The geometric margin of this solution turns out to be negative (0.75-0.75). For this setup, we use a network with LR: 0.0001; Embedding Init: PyTorch-Default ($\mathcal{N}(0,1)$), Weight Decay: Off, Dropout: Off
(a) PCA
(b) Emb. Heatmap vs. Adjacency
(c) Eigenvector Projections
Figure 29: Higher learning rate results in more geometry: Increasing the learning rate from $10^{-4}$ in Fig.˜28 to $10^{-1}$ results in a more geometric solution. For this setup, we use a network with LR: 0.1; Embedding Init: PyTorch-Default ($\mathcal{N}(0,1)$), Weight Decay: Off, Dropout: Off
(a) PCA
(b) Emb. Heatmap vs. Adjacency
(c) Eigenvector Projections
Figure 30: Higher weight decay results in more geometry: Adding weight decay to the setting of Fig.˜28 results in a more geometric (although zigzagging) solution. For this setup, we use a network with LR: 0.0001; Embedding Init: PyTorch-Default ($\mathcal{N}(0,1)$), Weight Decay: 10, Dropout: Off
(a) PCA
(b) Emb. Heatmap vs. Adjacency
(c) Eigenvector Projections
Figure 31: Higher dropout results in more geometry: Adding dropout to the setting of Fig.˜28 and a slight increase in the learning rate results in a more geometric (although zigzagging) solution. For this setup, we use LR: 0.001; Embedding Init: PyTorch-Default ($\mathcal{N}(0,1)$), Weight Decay: Off, Dropout: 0.4
(a) PCA Embeddings
(b) Embedding Heatmap and Adjacency Matrix
(c) Eigenvector Projections
Figure 32: Unidirectional graph can still result in geometry: At least on small graphs, and for the Transformer (TinyGPT), training on only one direction of the edges suffices for the model to store geometrically. Similar results hold for our other graphs.
@@ Appendix D Edge supervision and training dynamics
There are tangential aspects of our large path-star graph training that are worth elaborating on: the roles of reverse edges (§D.1), of pause tokens (§D.2), of interleaving edge-memorization (§D.3).
@@ D.1 The role of reverse edges
Reverse edges seem to play a nuanced role in our observations. On the large path-star task in §2, we find it necessary to augment training on the reverse edges; on the other hand, for the tiny graphs, a geometry arises even without these reverse edges. Perhaps, reverse edges are needed for larger tasks; or perhaps, they are necessary to perform implicit reasoning and retrieval. We leave it for future work to gain greater clarity on this effect, which is tied to the reversal curse [Berglund et al., 2024, Allen-Zhu and Li, 2023].
@@ D.1.1 The critical role of reverse edges in the large path-star task
(Role of reverse edges) We find in Fig.˜33 that:
@@
    A Transformer trained on only the forward edges, struggles on both forward and reverse path-finding tasks (see the middle color in Fig.˜33).
@@
    A Transformer trained on only the reverse edges, achieves non-trivial accuracy on the reverse path-finding task; however, it fails on the forward path-finding task (see the third color in Fig.˜33).
@@
We note that the asymmetry between forward and reversed path-generation tasks stems from their algorithmic complexity. The reversed path generation is algorithmically trivial on path-star graphs because each node has a unique predecessor along any target path—the model simply needs to follow the unique backward edges. Forward generation, however, requires planning (examine each outgoing path) or lookahead (track the reverse path without explicit chain-of-thought and reverse it). Indeed, for the in-context task of B&N’24, the model fails on the forward task, but strikingly succeeds on the reverse task, as corroborated in Fig.˜34 (right). Even in the in-weights setting Fig.˜34 (left), the reverse path is generally quicker to learn and yields higher accuracy.
@@ D.2 Pause tokens for computational slack
In the same in-weights path-finding task of §2, we find that it is helpful to insert pause tokens [Burtsev et al., 2020, Goyal et al., 2024] to achieve quicker accuracy gains during training. Pause tokens are added by appending dummy tokens to the prefix of the path-finding task both during training and inference.
Fig.˜35 shows that adding a short sequence of pause tokens after the prompt reliably boosts exact-match accuracy across graphs, for a given amount of training time. Increasing the number of pause tokens increases speed of convergence.
@@
In all our experiments, we have interleaved edge-memorization examples with path-finding examples. An alternative training method would be a two-phased approach, where we first enforce edge-memorization, and then follow up by finetuning on the path-finding task. We found this to be less stable, e.g., the model achieves a peak accuracy momentarily, only to deteriorate dramatically right after. This is a manifestation of the well-known effect that finetuning has on parametric memory [Li et al., 2024, Luo et al., 2025]. Since this is a confounding effect, we do not choose this regime for our experiments.
However, we confirm that even in this regime, our models do achieve a high peak accuracy (see Fig.˜36). The fact that the composition task is learnable in this regime implies that the edge-pretrained model must have come with an adequate global geometry despite being trained only on local supervision (Section˜3.1).
@@ E.1 Empirical failure of composition learning under associative memory
As discussed in §2.3, it is well-known that certain compositional tasks are hard to learn empirically; theoretically, this has been proven in a certain sense. However, these prior discussions involve composing information available in the context, rather than in the weights. To test this intuition in our in-weights composition task, we design an experiment where we freeze the embeddings of a Transformer and train it on our path-star task. If the model succeeded in this task, it may mean one of two things: either the model develops a geometric memory in a subsequent layer, or the model does in fact efficiently learn how to compose associative matrix operations—going against our intuition derived from prior limits on composition learning. However, we find that even after 50,00050{,}000 training epochs (using the same GPT-mid architecture described in Table 2 and §B.2.1, except with frozen token embeddings; and the optimization hyperparameter grid search reported in §B.3), the model fails to learn the in-weights path-star task. We believe, this is preliminary evidence that associative memory indeed struggles to learn compositional in-weights tasks. A fleshed-out proof of this negative result is left for future work.
@@
In datasets where redundancies exist, the complexity of a lookup table scales quickly with the training set size (say $n$), whereas the more succinct solution does not (or at worst, grows polynomially slower). For example, if the data is linearly separable in some constant dimensionality, the linear classifier can be described in $\Theta(1)$ bits, whereas a lookup table (such as a nearest neighbor model) would require $\Theta(n)$ bits. However, this wide disparity in complexity does not necessarily surface in our setting, which is a memorization task without redundancies. Concretely, at least in terms of bits and $\ell_{2}$ norms, there are simple graphs where both the geometric and associative views are equally complex— and at worst, differ by a constant factor of at most $2$ or so depending on whether edge/weight symmetries are constrained during training on not. Thus, the gap does not scale as $\Theta(n)$, which in our case would be the edge count or the vertex count of our graph.
We make two notes about the constant multiplicative factor of $2$ in the gap between the two complexities. First, the factor does not appear if only direction of the edges are stored, or if the embedding weights are untied. Second, it may seem that this constant factor may partly explain the geometric bias; however, a geometry appears even when memorizing only one direction of the edges in the smaller graphs (see Fig.˜32), where both forms of storage must be equally succinct.
Proof intuition. We first roughly derive the bit and norm complexity for a general graph. The key idea is that associative memory scales with the edge count, whereas a geometric memory scales with the vertex count. Then, we argue how this resolves to similar values for graphs like the path-star or a cycle, where the edge count and the vertex count are the same (plus or minus $1$).
@@
In the standard associative memory view that we have discussed, associations are represented through the function ${\bf{\Phi}}(v)^{T}{\bf{W}}_{\tt assoc}{\bf{\Phi}}(u)$. However, weight-tied dual encoder models like Node2Vec can only represent functions of the form ${\bf{\Phi}}(v)^{T}{\bf{\Phi}}(u)$. When weight-tied, this precludes the form of associative memory we care about; but it still allows two forms of associative memories under special conditions: if the graph is bipartite (Proposition˜3) or if the embedding dimensionality is sufficiently large (in which case, the construction is contrived; see Proposition˜5). We can also show that non-weight-tied models can represent any graph associatively (see Proposition˜4). In all these results below, we will show that our constructions are associative in that they achieve a logit of $1$ on adjacent vertices, but a logit of 0 on any non-adjacent vertices. In this sense, only local associations are captured, whereas no global geometry is.
@@
For each node, we take its corresponding row from the $|{V}|\times|E|$ incidence matrix. In other words, the embedding is $|E|$-dimensional, where the $i$th dimension corresponds to the $i$th edge; for edge $(u,v)$, the embedding of $u$ and $v$ both are set to $1$ along the dimension corresponding to $(u,v)$. Notationally, if the edges are indexed as $1,2,\ldots$, then ${\bf{\Phi}}(u)_{i}=\mathbf{1}[u\in e_{i}]$, where $\bf{1}$ is the indicator function. Then, we have that ${\bf{\Phi}}(u)\cdot{\bf{\Phi}}(v)=\mathbf{1}[(u,v)\in E]$. Thus, the dot products capture only local information. ∎